If the length of the shorter side of a right triangle is 11 cm and the length of the hypotenuse is 61 cm, what is the length of the longer side?
Answer by checkley79(3341)

A^2+B^2=C^2
11^2+B^2=61^2
121+B^2=3,721
B^2=1,371-121
B^2=3,600
B=SQRT3,600
B=60 ANS. FOR THE LONGER SIDE.
PROOF:
11^2+60^2=81^2
121+3,600=3.721
3,721=3,721